OpenStudy (anonymous):

4 8/15 how to make this into a decimal

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

fast please!

OpenStudy (amistre64):

4 8/15 = 4 + 8/15 4 + 8/15 = \(\Large \frac1{15}\)(4(15)+8) ....fast?? i assume your taking a timed test then by that comment ....

OpenStudy (anonymous):

no because my test is tomorrow and it lmost dinner sry

OpenStudy (anonymous):

so the answer is 4.5333333333333?

OpenStudy (amistre64):

4(15) + 8 = 68 therefore just divide 68 by 15, longhand or with a calculuator

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

thanks!

OpenStudy (amistre64):

yep ;)

OpenStudy (anonymous):

another question